OpenStudy (anonymous):

what is the domain of f(x)=|-x+6|+2

OpenStudy (amistre64):

it doesnt seem to have any need for restrictions

OpenStudy (anonymous):

\[That's \right. the domain: x \in \mathbb{R}\]. all real numbers.

OpenStudy (amistre64):

right is a special word when it has the \ in front of it

OpenStudy (amistre64):

youd have to space that out

OpenStudy (amistre64):

that's\ right.\ the\ domain .. as such, or use the \text{} option

OpenStudy (amistre64):

also, i dont think they use the \mathbb tag ... \[\mathbb{R}\] i huess it does

OpenStudy (anonymous):

super, duper! thanks.